OpenStudy (anonymous):

a culture started with 6000 bacteria. after 6hrs, it grew to 7200 bacteria. predict how many bacteria will present in 17 hrs. does anybody know the answer its my last question.. /:

OpenStudy (anonymous):

It grew 1200 bacteria in 6 hours. This means 200 an hour.

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

No, calculate it for 17 hours and add the 6000 from the beginning.

Parth (parthkohli):

No, the answer is\[7200 + 200\cdot17\]

OpenStudy (anonymous):

10600 is the answer?0:

OpenStudy (anonymous):

Actually, I think it's already 6 hours in. so you just add 12 hours? I find this question could have been better explained to avoid any errors.

OpenStudy (anonymous):

AHhhh never mind youère right.

hartnn (hartnn):

@ParthKohli , no it isn't.

OpenStudy (anonymous):

17 hours from there, sorry.

Parth (parthkohli):

Right, that could be a case as well.\[6000 + 200\cdot 17\]

OpenStudy (anonymous):

But it says will present in 17 hours.

OpenStudy (anonymous):

wait.... dang so confusing.

OpenStudy (anonymous):

yepp,

Parth (parthkohli):

The answer is one of the following:\[1) \quad 6000 + 200\cdot 17\]\[2) \quad 7200 + 200\cdot 17\]

OpenStudy (unklerhaukus):

\[B(t)=B_02^{t/{T}}\]

OpenStudy (anonymous):

I'd write both honestly. On a test there would never be such a question anyways.

OpenStudy (anonymous):

so much fail.

hartnn (hartnn):

@dom_14 u have options/choices ?

OpenStudy (anonymous):

no its just type in the answer , and if i get it wrong then i fail.

OpenStudy (anonymous):

Wow...

hartnn (hartnn):

@UnkleRhaukus i think its simple math question with linear growth rate. so in 17 hours it will be 6000+200*17 =9400

OpenStudy (anonymous):

Most likely.

OpenStudy (unklerhaukus):

\[B(t)=B_02^{t/{T}}\] \[7200=6000\cdot2^{6/{T}}\] \[\frac{7200}{6000}=2^{6/T}\] \[\log_2\left(\frac{7200}{6000}\right)=6/T\] \[T=\frac{6}{\log_2\left(\frac{7200}{6000}\right)}\]

OpenStudy (unklerhaukus):

Bactria grow exponentially

OpenStudy (unklerhaukus):

find T the doubling time , then substitute into the first equation

OpenStudy (anonymous):

does it make sense @dom_14 ?

OpenStudy (anonymous):

thanks guys.

OpenStudy (anonymous):

@dom_14 did you pass? don't leave us hanging in suspense...

OpenStudy (anonymous):

no i needed 8 to pass and i only got 7 right.. darn.

OpenStudy (anonymous):

It was pretty hard.

OpenStudy (anonymous):

I want to do that test :/

OpenStudy (anonymous):

yeah, it is. like im so tired of taking it and no you dnt, it switches up questions.

OpenStudy (anonymous):

What's it called?

OpenStudy (anonymous):

you'll have to update your FB status I guess :(

OpenStudy (anonymous):

plato.